simplify {12^1/5 / 27^1/5}^5/2
please solve tis it is important

Grade:9

1 Answers

Ram Kushwah
3 years ago
{12^1/5 / 27^1/5 } ^5/2
conmbining the numbers of numerator and denominator
={(12/27)^1/5}^5/2
=(12/27)^(1/5*5/2)
=(12/27)^1/2
=(4/9)^1/2
=2/3 ans